# The False Promise of Zero-Shot Super- Resolution in Machine-Learned Operators

### v1

@@ 4.1 Physical Optimization Constraints
Physics-informed optimization constraints have been proposed as a means of achieving accurate inference in the zero-shot super-resolution setting (Li et al., 2024b). For each dataset in {Darcy, Burgers, Navier Stokes}, we train FNOs at all avaliable resolutions. We optimize each set of model parameters $\theta$ with a dual optimization objective $\mathcal{L}(\theta)=(1-w)*\ell_{\text{data}}(\theta)+w*\ell_{\text{phys}}(\theta)$, where $\ell_{\text{data}}$ is the original data-driven loss (mean squared error), and where $\ell_{\text{phys}}$ is an additional physics-informed loss, which explicitly enforces that the governing PDE is satisfied. We use the physics losses of Li et al. (2024b), and we detail their implementation in Appendix E.
We find that the data-driven loss always outperforms any training objective that includes a physics constraint (see Appendix E: Fig. 17). We determine this by evaluating $w\in$ {0, 0.1, 0.25, 0.5} for {Darcy, Burgers, Navier Stokes} at resolutions 64, 512, 255, respectively. Among the objectives that contain a physics constraint, we observe a clear trend: the lower the physics constraint is weighted, the better test performance the model achieves. This indicates that physics constraints make it more challenging for the model to be trained optimally despite extensive hyperparameter tuning (perhaps due to well known practical reasons such as being difficult to optimize (Krishnapriyan et al., 2021; Subramanian et al., 2022; Gao & Wang, 2023; Wang et al., 2023)).
To further illustrate this, we use the smallest non-zero $w=0.1$, and compare the physics-informed optimization with solely data-driven optimization. In Fig. 6, the predicted spectra of data from models optimized with a physics loss generally diverge more substantially across test resolutions than models optimized with only a data loss. Models optimized with physics constraints even fail to accurately fit their training distributions (Fig. 6(c)), and they fail to generalize to both super- and sub-resolution data (Fig. 6(a,b,d)). See Appendix E for results on all datasets. We conclude that physics informed constraints do not reliably enable multi-resolution generalization.
Figure 6: (Physics-Informed) Optimization Evaluation. FNO trained on Burgers data at resolution 256 (indicated by *), with and without physics optimization constrains. Average test spectra are visualized. The spectra generated by the models trained with physics+data loss do not match the ground truth. (Physics Loss term weighting $w=$ 10%.) Full results are in Appendix E, Figs. 19-21.
@@ 4.2 Band-limited Learning
We study two approaches which propose learning band-limited representations of data: convolutional neural operators (CNO) (Raonic et al., 2023); and the Cross-Resolution Operator-Learning (CROP) pipeline (Gao et al., 2025). Both CNO and CROP have been proposed as alias-free solutions to enable multi-resolution inference (Bartolucci et al., 2023). CNOs are fixed-resolution models; to use them, one must first interpolate their input to the model’s training resolution, do a forward pass, and then interpolate back to the original dimension. The CROP pipeline is more general and extends to any class of model by interpolating inputs (both at training and inference) to and from a fixed-dimensional representation before and after a forward pass.
We train and evaluate a CNO and CROP+FNO model on all {Darcy, Burgers, Navier Stokes} datasets. Models are optimized using optimal hyper-parameters which were found via a grid search detailed in Appendix B. In Fig. 7, we visualize the predicted spectra of models trained on resolution 16 data across test data of resolution {16, 32, 64, 128}. We observe that the CNO does accurately learn the band-limited representation of its training data. That is, the spectra matches that of the ground truth until frequency 8ff, after which, by design, it drops sharply. This means that while CNO does not alias, it cannot predict frequencies higher than what was seen during training. Similarly, we observe that the CROP model, accurately fits lower frequencies, but struggles to fit high frequencies accurately across resolutions.
We note more broadly that the design of band-limiting a model’s training data and predictive capacity is counter-intuitive to the goal of multi-resolution inference, in which, a broad range of frequencies must be modeled accurately. Band-limiting a model’s predictive capacity may enable accurate fixed-resolution representations, but it ensures that high-frequency information is not predicted accurately (or at all). We conclude that band-limited learning limits a model’s utility for multi-resolution inference (full results are in Appendix F).
@@ 5 Multi-Resolution Training
We hypothesize that the reason models struggle to do zero-shot multi-resolution inference is because data representing a physical system at varying resolutions is sufficiently out-of-distribution to the model’s fixed-resolution training data. To remedy this, we propose a data-driven solution: multi-resolution training (i.e., training on more than one resolution).
@@
Figure 8: Multi-resolution training. FNO trained on multi-resolution Darcy data. Top row: Average number of pixels in a data sample in the training set. Lower number of pixels enables faster data generation and model training. Middle row: The ratios of data within each resolution bucket. Bottom row: The average test loss across different resolutions. Lower loss is better. Notice that mixed resolution datasets achieve both low average data size and low loss (ideal scenario).
@@
We compose multi-resolution datasets by randomly sampling different proportions of training data at varying resolutions $r_{1},...,r_{n}$}, where $r_{x}$ is the proportion of $x$ resolution training data, and where $n=4$. We begin by evaluating dual-resolution training. Li et al. (2024a) have previously shown that dual-resolution active learning enables more accurate high-resolution inference. We extend this finding and evaluate whether dual-resolution training can enable accurate multi-resolution inference. For each dataset in {Darcy, Burgers, Navier Stokes}, we combine data across resolutions in a pairwise manner creating $\frac{n(n-1)}{2}$ dual-resolution sets; the ratio of data samples between the two resolutions is varied over $p\in$ {0.05, 0.1, 0.25, 0.5, 0.75, 0.9, 0.95}. In Fig. 8($a$-$f$), we observe for pair-wise training, the test performance for data that corresponds to the two training resolutions is generally better, but we also note that there are not consistent gains for the two non-training resolutions. This indicates that models perform best on the data resolutions on which they are trained.
@@
To improve multi-resolution capabilities, we investigate the impact of including data from all resolutions. We first assess an equal number of samples across resolutions. In Fig. 8($g$), the test performance across all resolutions improves, which confirms that multi-resolution training benefits multi-resolution inference. Next, we ask: Can we improve the computationally efficiency of multi-resolution training? To do this, the training dataset must be composed of primarily low resolution data, as it is both the cheapest to generate and the cheapest to train over (see Appendix G: Fig. 31). We compose two additional multi-resolution datasets: {(0.7, 0.1, 0.1, 0.1), (0.9, 0.5, 0.3, 0.2)}. In Fig. 8($h,i$), models remain competitive across test resolutions, even as we decrease the amount of high-resolution data. In Fig. 9, we observe the consistent trend for all datasets: models are able to achieve a balance between dataset size and multi-resolution test loss via multi-resolution training. Optimizing the ratios across all resolutions remains an exciting future direction. Full results can be found in Appendix G.
@@ 6 Related Work
Modeling PDEs via Deep Learning. Three prominent approaches exist to discretely model PDEs with deep learning. 1) Mesh-Free models learn the solution operator to a specific instance of a PDE (Yu et al., 2018; Raissi et al., 2019; Bar & Sochen, 2019; Smith et al., 2020; Pan & Duraisamy, 2020). Mesh-free models can be queried to return a measurement at any time and space coordinate. While this means that a single model can resolve its solution at arbitrary discretization, this approach has two shortcomings: (i) Inference costs increases with number of points queried; and (ii) Models cannot generalize beyond the specific PDE instance it was trained on. 2) Fixed-Mesh models remedy both issues by learning a solution operator for a PDE family over a fixed-resolution mesh (Guo et al., 2016; Zhu & Zabaras, 2018; Adler & Öktem, 2017; Bhatnagar et al., 2019; Khoo et al., 2021). The inference costs of fixed-mesh models are lower than traditional numerical methods at corresponding mesh resolutions. However, fixed-mesh models fall short when one is interested in modeling scale phenomena that cannot be resolved via the fixed-mesh resolution (e.g., high-frequency information in turbulent systems). 3) Mesh-invariant models, unlike fixed-mesh models, are capable of doing inference at arbitrary mesh resolutions (Li et al., 2020b, a; Lu et al., 2021; Bhattacharya et al., 2021; Nelsen & Stuart, 2021; Patel et al., 2021; Rahman et al., 2022; Fanaskov & Oseledets, 2023). These methods have been proposed as a means to learn mesh-invariant solution operators to entire PDE families cheaply: train on low-resolution data, and use in a zero-shot fashion on arbitrary resolution data (e.g., zero-shot multi-resolution). In this work, we examine the zero-shot multi-resolution utility of mesh-invariant models.
@@
Aliasing (and corrective measures) in Deep Learning. Sources of aliasing in deep learning include both artifacts of a pixel grid (which models learn and amplify) and the application of point-wise non-nonlinearities to intermediate model representations (Karras et al., 2021; Gruver et al., 2022; Wilson, 2025). A straightforward approach, first introduced in generative adversarial networks, to stem nonlinearity-caused aliasing is to up sample a signal before applying a non-linearity, followed by down sampling the signal (Karras et al., 2021). Bartolucci et al. (2023) and Raonic et al. (2023) extend the application of an anti-alias activation function design to scientific machine learning. This does prevent aliasing, but it does not enable models trained at a specific resolution to resolve higher frequencies in higher resolution data. Gao et al. (2025) proposed a framework that “lifts” arbitrarily discretization data to a fixed-resolution band-limited space, both to train a model and to perform inference. We investigate the sources of aliasing the context of zero-shot multi-resolution inference, and we show that proposed solutions fall short in remedying the core issue: challenges with out-of-distribution generalization.
@@ 7 Conclusion and Future Work
For MLOs to be as versatile as numerical methods-based approaches for modeling PDEs, they must be able to perform accurate multi-resolution inference. To better understand an MLO’s abilities, we break down the task of multi-resolution inference by assessing a trained model’s ability both to extrapolate to higher/lower frequency information in data and to interpolate across varying data resolutions. We find that models trained on low resolution data and used for inference on high-resolution data can neither extrapolate nor interpolate; and, more generally, they fail to achieve accurate multi-resolution inference. Changing the resolution of data at inference time is akin to out-of-distribution inference: models have not learned how to generalize in such settings. We document that models used in a zero-shot multi-resolution setting are prone to aliasing. We study the utility of two existing solutions–physics-informed constrains and learning band-limited learning–and find that neither enable accurate multi-resolution inference.
Based on these results, we introduce a simple, intuitive, and principled approach to enable accurate multi-resolution inference: multi-resolution training. We show that models perform best at resolutions on which they have been trained; and we demonstrate that one can computationally efficiently achieve the benefits of multi-resolution training via datasets composed with mostly low-resolution data and small amounts of high-resolution data. This enables accurate multi-resolution learning, with the added benefit of low data-generation and model training cost. A promising future direction remains the automated selection of multi-resolution training data using strategies like active learning.
@@ Acknowledgements
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number DE-SC0023112.
@@ Appendix B Hyper-parameter Search
Table 1: Optimal FNO hyper-parameters from hyper-parameter search outlined in Appendix B. *NS batch size had to be reduced to 1 for multi-resolution training experiments (see Sec. 5), and therefore we used a lower learning rate in that setting. ww=Physics Loss Coefficient (see Sec. 4.1).
@@
Table 2: Optimal CNO/CROP hyper-parameters from hyper-parameter search outlined in Appendix B. *The original CROP implementation did not include a 1D version, so we omit CROP for the 1D Burgers dataset.
@@ Appendix G Multi-Resolution Training
Appendix H Model Implementations
The Fourier Neural Operator is described in detail in (Li et al., 2020a); we closely follow their implementation which can be found at https://neuraloperator.github.io/dev/index.html.
The CNO is described in detail in (Raonic et al., 2023); we closely follow their implementation which can be found at https://github.com/camlab-ethz/ConvolutionalNeuralOperator/tree/main.
The CROP pipeline is described in detail in (Gao et al., 2025); we closely follow their implementation for CROP+FNO which can be found at https://github.com/wenhangao21/ICLR25-CROP/tree/main. We note that they did not include a 1D CROP implementation, so we exclude evaluation of CROP on the 1D Burgers dataset.

### v2

@@ 4.1 Physical Optimization Constraints
Physics-informed optimization constraints have been proposed as a means of achieving accurate inference in the zero-shot super-resolution setting (Li et al., 2024b). For each dataset in {Darcy, Burgers, Navier Stokes}, we train FNOs at all avaliable resolutions. We optimize each set of model parameters $\theta$ with a dual optimization objective $\mathcal{L}(\theta)=(1-w)*\ell_{\text{data}}(\theta)+w*\ell_{\text{phys}}(\theta)$, where $\ell_{\text{data}}$ is the original data-driven loss (mean squared error) with an additional physics-informed loss $\ell_{\text{phys}}$, which explicitly enforces that the governing partial differential equation is satisfied. We use the physics losses of Li et al. (2024b) and detail the implementation in Appendix E.
We find that the data-driven loss always outperforms any training objective that includes a physics constraint (Fig. 17). We determine this by evaluating $w\in$ {0, 0.1, 0.25, 0.5} for {Darcy, Burgers, Navier Stokes} at resolutions 64, 512, 255, respectively. Among the objectives that contain a physics constraint, we observe a clear trend: the lower the physics constraint is weighted, the better test performance the model achieves. This indicates that physics constraints make it more challenging for the model to be trained optimally despite extensive hyperparameter tuning (perhaps due to practical reasons such as being difficult to optimize (Krishnapriyan et al., 2021; Subramanian et al., 2022; Gao and Wang, 2023; Wang et al., 2023)).
To further illustrate, we use the smallest $w=0.1$ and compare the physics-informed optimization with solely data-driven optimization. In Fig. 6, the predicted spectra of data from models optimized with physics loss generally diverge more substantially across test resolutions than models optimized with only a data loss. Models optimized with physics constraints even fail to accurately fit their training distributions (Fig. 6(c)), and fail to generalize to both super- and sub-resolution data (Fig. 6(a,b,d)). See Appendix E for results on all datasets. We conclude that physics informed constraints do not reliably enable multi-resolution generalization.
Figure 6: (Physics-Informed) Optimization Evaluation. FNO trained on Burgers data at resolution 256 (indicated by *) with and without physics optimization constrains. Average test spectra visualized. Spectra generated by the models trained with physics+data loss do not match the ground truth. Physics Loss term weighting $w=$ 10%. Full results in  Appendix E, Figs. 19-21.
@@ 4.2 Band-limited Learning
We study two approaches which propose learning band-limited representations of data: convolutional neural operators (CNO) (Raonic et al., 2023) and the Cross-Resolution Operator-Learning (CROP) pipeline (Gao et al., 2025). Both CNO and CROP have been proposed as alias-free solutions to enable multi-resolution inference (Bartolucci et al., 2023). CNOs are fixed-resolution models; to use them, one must first interpolate their input to the model’s training resolution, do a forward pass, and then interpolate back to the original dimension. The CROP pipeline is more general and extends to any class of model by interpolating inputs (both at training and inference) to and from a fixed-dimensional representation before and after a forward pass.
We train and evaluate a CNO and CROP+FNO model on all {Darcy, Burgers, Navier Stokes} datasets. Models are optimized using optimal hyper-parameters which were found via a grid search detailed in Appendix B. In Fig. 7, we visualize the predicted spectra of model’s trained on resolution 16 data across test data of resolution {16, 32, 64, 128}. We observe that the CNO does accurately learn the band-limited representation of its training data: the spectra matches that of the ground truth until frequency 8ff after which, by design, it drops sharply. This means that while CNO does not alias, it cannot predict frequencies higher than what was seen during training. Similarly, we observe that the CROP model, accurately fits lower frequencies, but struggles to fit high frequencies accurately across resolutions.
We note more broadly that the design of band-limiting a model’s training data and predictive capacity is counter-intuitive to the goal of multi-resolution inference, in which, a broad range of frequencies must be modeled accurately. Band-limiting a model’s predictive capacity may enable accurate fixed-resolution representations, but ensures that high-frequency information is not predicted accurately (or at all). We conclude that band-limited learning limits a model’s utility for multi-resolution inference (full results in Appendix F).
@@ 5 Multi-Resolution Training
We hypothesize that the reason models struggle to do zero-shot multi-resolution inference is because data representing a physical system at varying resolutions is sufficiently out-of-distribution to a model’s fixed-resolution training data. To remedy this, we propose a data-driven solution: multi-resolution training (i.e., training on more than one resolution) (Li et al., 2024a; Rowbottom et al., 2025; Berner et al., 2025; George et al., 2024; Tang et al., 2024; Lyu et al., 2023; Ma et al., 2024; Lanthaler et al., 2024).
@@
Figure 8: Multi-resolution training. FNO trained on multi-resolution Darcy data. Top row: Average number of pixels in a data sample in the training set. Lower number of pixels enables faster data generation and model training. Middle row: The ratios of data within each resolution bucket. Bottom row: The average test loss across different resolutions. Lower loss is better. Notice: mixed resolution datasets achieve both low average data size and low loss (ideal scenario).
@@
We compose multi-resolution datasets by randomly sampling different proportions of training data at varying resolutions $r_{1},...,r_{n}$} where $r_{x}$ is the proportion of $x$ resolution training data and $n=4$. We begin by evaluating dual-resolution training. Li et al. (2024a) have previously shown that dual-resolution active learning enables more accurate high-resolution inference. We extend this finding and evaluate if dual-resolution training can enable accurate multi-resolution inference. For each dataset in {Darcy, Burgers, Navier Stokes}, we combine data across resolutions in a pairwise manner creating $\frac{n(n-1)}{2}$ dual-resolution sets; the ratio of data samples between the two resolutions is varied over $p\in$ {0.5,0.1,0.25,0.5,0.75,0.9,0.1}. In Fig. 8($a$-$f$), we observe for pair-wise training, the test performance for data that corresponds to the two training resolutions is generally better, but there are not consistent gains for the two non-training resolutions. This indicates that models perform best on the data resolutions on which they are trained.
@@
To improve multi-resolution capabilities, we investigate the impact of including data from all resolutions. We first assess an equal number of samples across resolutions. In Fig. 8($g$), the test performance across all resolutions improves which confirms that multi-resolution training benefits multi-resolution inference. Next, we ask: Can we improve the computationally efficiency of multi-resolution training? To do this, the training dataset must be composed of primarily low resolution data as it is both the cheapest to generate and train over (Fig. 32). We compose two additional multi-resolution datasets: {(0.7, 0.1, 0.1, 0.1), (0.9, 0.5, 0.3, 0.2)}. In Fig. 8($h,i$), models remain competitive across test resolutions, even as we decrease the amount of high-resolution data. In Fig. 9, we observe the consistent trend for all datasets: models are able to achieve a balance between dataset size and multi-resolution test loss via multi-resolution training. Optimizing the ratios across all resolutions remains an exciting future direction. Full results in Appendix G.
@@ 6 Related Work
Modeling PDEs via Deep Learning. Three prominent approaches exist to discretely model PDEs with deep learning. 1) Mesh-Free models learn the solution operator to a specific instance of a PDE (Yu and others, 2018; Raissi et al., 2019; Bar and Sochen, 2019; Smith et al., 2020; Pan and Duraisamy, 2020). Mesh-free models can be queried to return a measurement at any time and space coordinate. While, this approach means that a single model can resolve its solution at arbitrary discretization, it has two shortcomings: (i) Inference costs increases with number of points queried. (ii) Models cannot generalize beyond the specific PDE instance it was trained on. 2) Fixed-Mesh models remedy both issues by learning a solution operator for a PDE family over a fixed-resolution mesh (Guo et al., 2016; Zhu and Zabaras, 2018; Adler and Öktem, 2017; Bhatnagar et al., 2019; Khoo et al., 2021). The inference costs of fixed-mesh models are lower than traditional numerical methods at corresponding mesh resolutions. However, fixed-mesh models fall short when one is interested in modeling scale phenomena that cannot be resolved via the fixed-mesh resolution (e.g., high-frequency information in turbulent systems ). 3) Mesh-invariant models, unlike fixed-mesh models, are capable of doing inference at arbitrary mesh resolutions (Li et al., 2020b; a; Lu et al., 2021b; Bhattacharya et al., 2021; Nelsen and Stuart, 2021; Patel et al., 2021; Rahman et al., 2022; Fanaskov and Oseledets, 2023). They have been proposed as a means to learn mesh-invariant solution operators to entire PDE families cheaply: train on low-resolution data, and use in a zero-shot fashion on arbitrary resolution data (e.g., zero-shot multi-resolution). In this work, we examine the zero-shot multi-resolution utility of mesh-invariant models.
@@
Aliasing (and corrective measures) in Deep Learning. Sources of aliasing in deep learning include both artifacts of a pixel grid which models learn and amplify and the application of point-wise non-nonlinearities to intermediate model representations (Karras et al., 2021; Gruver et al., 2022; Wilson, 2025). A straightforward approach, first introduced in generative adversarial networks, to stem nonlinearity-caused aliasing is to up sample a signal before applying a non-linearity followed by down sampling the signal (Karras et al., 2021). Bartolucci et al. (2023) and Raonic et al. (2023) extend the application of anti-alias activation function design to scientific machine learning; while this does prevent aliasing, it does not enable models trained at a specific resolution to resolve higher frequencies in higher resolution data. Gao et al. (2025) proposed a framework that “lifts” arbitrarily discretization data to a fixed-resolution band-limited space, to both train and do inference in. We investigate the sources of aliasing the context of zero-shot multi-resolution inference and show that proposed solutions fall short in remedying the core issue: out-of-distribution generalization.
@@ 7 Conclusion and Future Work
For machine-learned operators to be as versatile as numerical methods-based approaches for modeling PDE’s they must perform accurate multi-resolution inference. To better understand an MLO’s abilities, we break down the task of multi-resolution inference and assess a trained model’s ability to both extrapolate to higher/lower frequency information in data and interpolate across varying data resolutions. We find that models trained on low resolution data and used for inference on high-resolution data can neither extrapolate nor interpolate, and therefore, more generally fail to do accurate multi-resolution inference. Changing the resolution of data at inference time is akin to out-of-distribution inference: models have not learned how to generalize in such settings. We document that models used in a zero-shot multi-resolution setting are prone to aliasing. We study the utility of two existing solutions–physics-informed constrains and learning band-limited learning–and find that neither enable accurate multi-resolution inference.
We introduce a simple and principled approach to enable accurate multi-resolution inference: multi-resolution training. We first show that models perform best at resolutions they have been trained on. We then extend this finding and demonstrate that one can computationally efficiently achieve the benefits of multi-resolution training via datasets composed with mostly low-resolution data and small amounts of high-resolution data. This enables accurate multi-resolution learning with the added benefit of low data-generation and model training cost. A promising future direction remains the automated selection of multi-resolution training data using strategies like active learning.
@@ Acknowledgements
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number DE-SC0023112. YY acknowledges support from DOE Award DE-SC0025584 and Dartmouth College. AT was funded by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC 2075 – 390740016 and acknowledges the support by the Stuttgart Center for Simulation Science (SimTech). We would also like to acknowledge the DOE Competitive Portfolios grant and the DOE SciGPT grant.
Reproducibility Statement
To enable reproduction of our results we will link to the code base upon acceptance. We additionally include details about hyperparameter tuning in Appendix B, models (and their configuration details) in Appendix H, datasets (and their configuration details) in Appendix A, data extrapolation/interpolation implementation details in Appendix C.1. To ensure reliability of results we include an additional sensitivity analysis in Appendix L. Unless otherwise specified, all experiments were run using Python 3.10 on an NVIDIA A100-PCIE-40GB with driver version 550.163.01; CUDA Toolkit version was 12.4.131; PyTorch version was 2.9.1+cu126; the “neuraloperator” library (Kossaifi et al., 2025) version was 2.0.0. We report results from replicating a subset of our experiments across multiple computing architectures to ensure reproducibility in Appendix N.
@@ Appendix B Hyper-parameter Search
Table 1: Optimal FNO hyper-parameters from hyper-parameter search outlined in Appendix B. *NS batch size had to be reduced to 1 for multi-resolution training experiments (see Sec. 5), therefore we used a lower learning rate in that setting. ww=Physic Loss Coefficient (see Sec. 4.1).
@@
Table 2: Optimal CNO/CROP/DeepONet hyper-parameters from hyper-parameter search outlined in Appendix B. *The original CROP implementation did not include a 1D version, so we omit CROP for the 1D Burgers dataset.
@@ Appendix C Information Extrapolation and Resolution Interpolation
C.1 Extrapolation and Interpolation Implementation Details
Down sampling Procedure: For the resolution interpolation experiments, data was sampled at different resolution by employing standard down sampling of the data (keeping every Nth element) from the maximal data resolution. This is the same down sampling procedure used by PDEBench (Takamoto et al., 2022).
Low-pass Filtering Procedure: For the information extrapolation experiments, data was filtered with a standard stop-band filter which restricts any frequency components above the desired filter limit $N$. Concretely, to low pass filter data, we transform the data to the frequency domain via an FFT and shift it such that the lowest frequency component is the center of the spectrum. We determine which frequency components are greater than our limit $N$ by assessing which components fall outside a radius of $N$ from the center; we then set the energy of any frequency components greater than $N$ to zero. Finally, the (filtered) data is transformed back to the spatial domain via an inverse FFT.
C.2 Full Results
@@ Appendix G Multi-Resolution Training
G.1 Multi-Resolution Training Performance Gains
Here we profile both the memory savings (see Tab. 3) and wall-clock training time (see Tab. 4) saving due to the reduction in dataset size from multi-resolution training compared with the single-resolution training counterpart. We empirically observe the smaller dataset size is associated with reduction in training time in Fig. 32.
Table 3: Impact of Multi-Resolution Training on Memory. Reporting the size of datasets in Gigabytes (GB) for multi-resolution dataset with ratios (0.9, 0.5, 0.3, 0.2) vs. single-resolution maximum resolution dataset along with the percent decrease in dataset size. The smaller the dataset, the better.
Dataset Single-Res. Size (GB) Multi-Res. Size (GB) % $(\downarrow)$
Darcy 0.54 0.02 96% $\downarrow$
Burgers 0.34 0.005 98% $\downarrow$
Navier Stokes 0.96 0.04 96% $\downarrow$
Table 4: Impact of Multi-Resolution Training on Wall-clock Training Time. Reporting the training time (over 150 epochs) for multi-resolution dataset with ratios (0.9, 0.5, 0.3, 0.2) vs. single-resolution maximum resolution dataset along with the percent decrease in dataset size. The lower the training time, the better.
Dataset Single-Res. Training Time (hours) Multi-Res. Training Time (hours) % $(\downarrow)$
Darcy 0.33 0.05 86% $\downarrow$
Burgers 6.03 5.8 3.2% $\downarrow$
Navier Stokes 21.9 9.6 56% $\downarrow$
Figure 32: Data size corresponds with training time. We notice a clear trend with Darcy and Navier Stokes datasets: as data size increases, so does average training time per epoch. For Burgers, this trend is less clear; however, we note that the Burgers dataset is several orders of magnitude smaller and therefore can be used with a high batch size (see Tab. 1) which reduces the computational gains achieved from a smaller sized dataset.
Appendix H Model Implementations and Configurations
The Fourier Neural Operator is described in detail in (Li et al., 2020a); we closely follow their implementation which can be found at https://neuraloperator.github.io/dev/index.html. We detail key model configuration parameter choices here:
  1. 1.
    Max Modes: When training and evaluating models to do zero-shot super-resolution or sub-resolution and multi-resolution training/inference, we train with the max modes parameter set to half the maximum training resolution. Further, we do an ablation study of the max modes parameter in Appendix I to assess impact of FNO’s ability to do information extrapolation and resolution interpolation.
  2. 2.
    Layer count: 4 (standard implementation)
  3. 3.
    Hidden channels: 32 (standard implementation)
  4. 4.
    Activation Function: All main paper experiments utilize the Gelu nonlinearity (Hendrycks, 2016). Further, we assess the impact of utilizing anti-aliasing activation function on zero-shot super and sub-resolution, information extrapolation, and resolution interpolation in Appendix K (Karras et al., 2021).
  5. 5.
    All other parameters: All other FNO parameters were initialized to default values detailed in the standard neuraloperator Python package (kossaifi2024library).
The CNO is described in detail in (Raonic et al., 2023); we closely follow their implementation which can be found at https://github.com/camlab-ethz/ConvolutionalNeuralOperator/tree/main. We detail key model configuration parameter choices here:
  1. 1.
    Residual block count: 6 (standard implementation)
  2. 2.
    Latent size: 64
  3. 3.
    Input Resolution: Varries based on training data resolution.
  4. 4.
    All other parameters: All other CNO parameters were initialized to default values detailed in the official CNO implementation: https://github.com/camlab-ethz/ConvolutionalNeuralOperator/tree/main.
The CROP pipeline is described in detail in (Gao et al., 2025); we closely follow their implementation for CROP+FNO which can be found at https://github.com/wenhangao21/ICLR25-CROP/tree/main. We note that they did not include a 1D CROP implementation, so we exclude evaluation of CROP on the 1D Burgers dataset. We detail key model configuration parameter choices here:
  1. 1.
    Hidden channels: 32 (standard implementation)
  2. 2.
    Latent size: 64
  3. 3.
    Max Mode: 32 (half latent size)
  4. 4.
    All other parameters: All other CROP parameters were initialized to default values detailed in the official CROP implementation: https://github.com/wenhangao21/ICLR25-CROP/tree/main.
The DeepONet model is described in detail in (Lu et al., 2021a); we closely follow the implementation of (Gao et al., 2025) which can be found at https://github.com/wenhangao21/ICLR25-CROP/tree/main. We detail key model configuration parameter choices here:
  1. 1.
    Branch layer count: 4 (standard implementation)
  2. 2.
    Trunk layer count: 4 (standard implementation)
  3. 3.
    Branch layer width: 128 (standard implementation)
  4. 4.
    Trunk layer width: 128 (standard implementation)
  5. 5.
    All other parameters: All other DeepONet parameters were initialized to default values detailed in the official implementation of (Gao et al., 2025): https://github.com/wenhangao21/ICLR25-CROP/tree/main.